In a certain city, the daily consumption of electric power, X (in millions of kilowatt hours) is a random variable where X ~ Gamma (α, β) with a mean of 6 and a variance of 12. i) Find α and β. ii) Given a month has 30 days, how many kilowatt hours should we expect the city to consume during the month? 


1
Expert's answer
2022-02-22T11:52:55-0500

"i)"

The formula for the mean and variance of the Gamma distribution are given as,

"E(x)=\alpha\beta" and "var(x)=\alpha\beta^2".

Now,

"\alpha\beta=6" and "\alpha\beta^2=12\implies (\alpha\beta)\beta=12"

Since "\alpha\beta=6\implies 6\beta=12\implies \beta=2"

"\alpha\beta =6" but "\beta=2\implies 2\alpha=6\implies \alpha=3"

Therefore, the values of "\alpha" and "\beta" are 3 and 2 respectively.

"ii)"

For this part, the distribution for the "n=30" independent days will be a Gamma distribution with parameters "(\displaystyle\sum^{30}_{i=1}\alpha_i, \beta )". That is, "Gamma (\displaystyle\sum^{30}_{i=1}\alpha_i, \beta )=Gamma((60),2)" . Its mean is "60\times 2=120".

Therefore, we expect that 120 Kilowatt hours will be consumed by the city during this month .